OpenStudy (leslielopez11):

Solve the following system of equations by substitution and select the correct answer below: 2x − 4y = 24 4x − 12y = 36

OpenStudy (anglina4196):

you first have to choose which variable you will substitute

OpenStudy (ajspeller):

Take equation #1 and solve it for either x or y. Then take that expression for either x or y and substitute it into equation #2. Now, you have only 1 variable and are now able to solve the equation. Take that answer and then substitute it back into either of the original equations to solve for the other variable. Here is a tutorial that shows you how to solve this questions using the substitution method. https://www.youtube.com/watch?v=boZn9Kml8Is

OpenStudy (anglina4196):

ok so if your solving it with substitution method. You have to make one of the equations Slope-Intercept form. so...... 2x-4y=24 -2x -2x -4y=-2x+24 ----------- -4 y=2/4x-6

OpenStudy (anglina4196):

or ....... 2x-4y=24 +4y +4y 2x=4y+24 ---------- 2 x=2y+12 lets use this one and substitute it into the other problem

OpenStudy (anglina4196):